# Interpretable Graph Learning Over Sets of Temporally-Sparse Data

### v2

@@ 3.2 Node, graph and set importance
[…]
@@
[…]
For graphs that are mixed non-linearly, i.e., graphs that belong in subsets of size greater than one, interpretability is more limited, and we can only provide the total contribution of the set to the final prediction
[…]
@@ 3.3 Expressivity properties
In this section we provide a theoretical analysis of the expressivness of GMAN.
@@
GMAN is strictly more expressive than GNAN.
The following theorem shows that a GMAN which is applied to subsets of graphs of size at least two, is more expressive than a GMAN that is applied to only subsets of size one:
@@
Let $S$ be a set of graphs $\{G_{i}\}_{j=1}^{m}$. Let $S_{1}=\{S_{i}\}_{i=1}^{m}$ be a partition of $S$ such that $|S_{i}|=1$. Let $S_{2}=\{S_{i}\}_{i=1}^{k}$ such that there exists $k$ with $|S_{k}|>1$. with a subset partition $\{S_{i}\}_{i=1}^{k}$. Then a GMAN trained over $S_{2}$ is strictly more expressive than a GMAN trained over $S_{1}$.
@@ 4 Empirical evaluation
In this section, we evaluate GMAN on a diverse set of real-world tasks to assess its effectiveness across both domains and graph structures  [^1] Code implementation can be found at http://github.com/azerio/Graph-Mixing-Additive-Networks---GMAN. We include two biomedical datasets: ICU mortality prediction and Crohn’s Disease (CD) diagnosis, where inputs are temporally sparse biomarker trajectories. To test performance on structurally different data, we also evaluate GMAN on a social media dataset for fake news detection, which involves more complex, densely connected graphs. This range of tasks enables a rigorous assessment of GMAN’s ability to generalize across graph structures and application settings, highlighting both its flexibility and practical relevance. Additional details on the datasets, experimental setup, and hyperparameter configurations are provided in the Appendix.
@@ 4.1 Medical predictions
We evaluate GMAN on two high-impact clinical prediction tasks: in-hospital mortality among intensive care patients and onset of Crohn’s Disease. In both settings, each individual is represented as a set of time-stamped biomarker trajectories, where each trajectory forms a directed line graph with nodes corresponding to test results and edges encoding the time elapsed between measurements. This representation preserves the temporal structure of each biomarker independently while enabling joint reasoning across biomarkers during learning. Both tasks are framed as binary classification. Importantly, the outcomes we predict carry serious clinical implications. Missed detections can delay critical treatment, while false alarms may lead to unnecessary interventions. These results demonstrate the effectiveness of GMAN when applied to real-world clinical data and its potential for deployment in high-stakes decision-making environments.
4.1.1 In-hospital mortality prediction
@@
The PhysioNet2012 (P12) dataset, introduced by [17], contains records from 11,988 ICU patients, following the exclusion of 12 samples deemed inappropriate according to the criteria in Horn et al. (2020). For each patient, time series measurements from 36 physiological signals (excluding weight) were recorded over the initial 48 hours of ICU admission. Additionally, each patient has a static profile comprising 9 features, including demographic and clinical attributes such as age and gender. The dataset is labelled for a binary classification task: predicting in-hospital mortality.
Table 1: Evaluation of GMAN on predicting in-hospital mortality over the P12 dataset.
Methods AUROC
Transformer 65.1 ± 5.6
Trans-mean 66.8 ± 4.2
GRU-D 67.2 ± 3.6
SeFT 66.8 ± 0.8
mTAND 65.3 ± 1.7
IP-Net 72.5 ± 2.4
DGM2 71.2 ± 2.5
MTGNN 67.5 ± 3.1
RAINDROP 72.1 ± 1.3
GMAN 76.64 ± 1.2
@@
We compare GMAN to other $9$ baselines evaluated in [55], including: Transformer [43], Trans-mean, GRU-D [8], SeFT [22], mTAND [41], IP-Net[40], DGM2[48] and MTGNN [49]. We use the data splits as in [55], which are random splits of train (80%), validation (10%), and test (10%) set over the data. We conducted a grid search by training on the training set and evaluating on the validation set. We selected the best performing model over the validation set. We report the average AUROC score and standard-deviation of the selected configuration with $3$ seeds. As input subsets for the model, we put all biomarkers in separate subsets, except for $5$ biomarkers which are in the same subset. The complete subsets information is provided in the Appendix.
@@ Results
The results are presented in Table 1. GMAN achieves the highest AUROC, outperforming the next best baseline, IP-Net by 4.14 AUCROC points. Notably, while other approaches are more expressive black-box methods, GMAN achieves high performance while remaining fully interpretable.
4.1.2 Crohn’s disease prediction
Data
The Danish health registries are comprehensive, nationwide databases covering healthcare interactions for over 9.5 million individuals [32]. A key resource is the Registry of Laboratory Results for Research (RLRR), which has collected laboratory test results from hospitals and general practitioners since 2015 [3]. From this data, we constructed a cohort of 8,567 individuals later diagnosed with Crohn’s Disease (CD) and 8,567 age-matched controls from the same registry. For each person, we extracted temporal trajectories of 17 routinely measured biomarkers, reflecting key physiological processes such as immune response, inflammation, organ function, and nutritional status. The only exception is faecal calprotectin (F-Cal), a stool-based biomarker specifically used to detect intestinal inflammation in conditions like CD. The complete list and descriptions of these biomarkers are provided in the Appendix. The task is binary classification: predicting future CD onset from pre-diagnostic medical histories.
Setup
We randomly split the dataset into 80% training, 10% validation, and 10% test sets. GMAN is trained for up to 300 epochs, with the best model selected based on validation performance. We report the mean and standard deviation (std) of AUROC across three random seeds. We trained GMAN using 7 subsets of biomarkers of different cardinality, representing different physiological functions. Full details of the subsets are provided in the Appendix.
We compare GMAN against widely used baselines for Crohn’s Disease prediction, including Random Forest (RF) [21] and Gradient Boosting (GB) [14], which are strong performers on structured biomedical data [15, 45, 37]. We also benchmark against two Transformer-based architectures, a standard Transformer[43] and Trans-mean [54]. We chose these as they act as high-capacity sequence models, designed to capture complex, long-range temporal dependencies without imposing strong parametric assumptions on the data-generating process. Because these baselines do not operate on graph-structured inputs, we revert to the original tabular format, where each row represents all biomarker tests taken on a given date, along with static features such as sex. For classical models, each row is treated as an independent example, with biomarker tuples parsed into numerical and categorical features. Categorical features are one-hot encoded, and missing values are imputed using a zero-inflation strategy.
Table 2: Evaluation of GMAN on predicting CD onset from pre-diagnostic biomarker trajectories.
Methods AUROC
Transformer 66.17 ± 0.3
Trans-mean 70.78 ± 2.4
Random Forest 80.76 ± 0.2
Naive Bayes 67.59 ± 1.8
GB 81.70 ± 0.3
GMAN 83.64 ± 0.9
Results
We report the results of GMAN in Table 2. GMAN outperforms all baseline models, achieving a 1.9-point improvement in AUROC over the next-best method. While classical baselines such as GB and RF offer feature importance scores, GMAN provides a richer and more granular interpretability framework. In addition to feature-level insights, it supports node-level and graph-level attributions, allowing us to pinpoint which specific measurements, and which temporal patterns, most influenced the model’s prediction. These multi-scale interpretability capabilities are explored in detail in the following subsection. Additional discussion of related work on predicting Crohn’s Disease onset, using both simple blood-based biomarkers and more complex biochemical features such as multi-omics data, is provided in the Appendix.
@@ 4.1.3 Clinical insights through interpretability
Figure 2: Node-level attributions for two individuals from (a) the P12 ICU mortality and (b) CD onset datasets. Node size indicates each measurement’s contribution to the prediction. GMAN highlights clinically relevant signals, renal and cardiac markers in P12, inflammation biomarkers in CD, demonstrating fine-grained temporal interpretability. CD data is privacy-protected via noise and temporal shifting.
Beyond predictive performance, a central strength of GMAN is that interpretability is built into the model by design, rather than added post hoc. This allows for fine-grained, temporally resolved explanations that go beyond global feature importance, enabling users to understand how and when specific biomarkers influence the model’s predictions. In high-stakes domains such as healthcare, this level of transparency is crucial. Clinical decision-making often depends not only on the outcome of a prediction but on a clear understanding of the reasoning behind it. Models that can provide such insights are far more likely to be trusted, audited, and integrated into clinical workflows. To highlight the interpretability of GMAN, we conduct attribution analyses on both the CD and P12 datasets, examining how the model assigns importance across time and features.
@@
In the GMAN framework, nodes represent individual measurements within a biomarker trajectory. As such, highly influential nodes can highlight critical phases where specific measurements most strongly impact the prediction. We use Equation 3 to quantify node-level contributions across biomarkers in both the CD and P12 datasets. Figure 2 displays node-level importance across biomarker trajectories for two randomly selected individuals from each dataset, with node size reflecting the magnitude of each node’s contribution to the model’s prediction. In both clinical tasks, the model’s attributions appear consistent with established biomedical knowledge. For Crohn’s Disease prediction, GMAN highlights key inflammatory and immune markers, such as faecal F-Cal, platelets, and lymphocytes, as primary contributors. All of these are known to play central roles in disease onset [45]. In the P12 example, the model assigns high importance to markers of renal function, liver injury, cardiac stress, and metabolic imbalance, aligning well with clinical predictors of poor outcomes in intensive care settings.
@@
In addition to node-level attributions, GMAN provides subset-level importance scores, quantifying the contribution of entire biomarker groups to the model’s prediction. This enables flexible, system-level interpretability, allowing users to assess the collective influence of physiologically related biomarkers on the target variable. We conduct a subset-level attribution analysis on the CD dataset using two clinically motivated grouping strategies. In the first, we assess individual biomarkers to determine whether the model prioritizes features known to be associated with CD onset, allowing us to verify its alignment with established biomedical knowledge. In the second, we group biomarkers into clinically coherent subsets based on physiological function, such as immune response; inflammation; oxygen transport; and liver function; to examine whether the model captures system-level patterns consistent with disease progression. Full details on the clinical significance of these groupings are provided in the Appendix.
(a) Single-biomarker groups GMAN
(b) Best-performing GMAN groups
@@
Figure 3: Subset-level contribution curves for Crohn’s Disease prediction. Each curve shows how the model’s output changes as increasing noise is added to the latent representation of a biomarker group. (a) uses individual biomarkers, revealing strong effects from F-Cal, platelets, CRP and lymphocytes. (b) uses physiologically coherent groups, where inflammation markers (CRP, F-Cal) dominate the prediction.
For each grouping strategy, we train a separate instance of GMAN. To quantify the importance of each biomarker subset, we use Equation 4, measuring how the model’s prediction changes as increasing amounts of noise are added to the features of nodes within that subset. To introduce noise in a structured way, we apply PCA [1] to the feature vectors of all nodes corresponding to biomarkers in the group, and progressively perturb the input along the 1st principal component. This procedure is performed independently for each subset and allows us to assess the model’s sensitivity to perturbations in biologically meaningful groupings, offering insight into the relative predictive weight of each group. Notice that as the importance of a subset is the exact value it contributes to the prediction, measuring the change in the prediction is equivalent to measuring the change in importance.
Figure 3 presents the results of the subset-level attribution analysis. In the single-biomarker setting (Figure 3(a)), F-Cal, platelets, and lymphocytes show strong directional effects on model output. F-Cal and platelets are positively associated with CD risk, while lymphocytes have an inverse effect, findings that are both biologically grounded and consistent with prior work [45]. In the clinically-coherent group setting (Figure 3(b)), the inflammation group, comprised of CRP and F-Cal, emerges as the most influential, with a pronounced non-linear effect on predictions, aligning with their established diagnostic relevance in CD [45].
@@ 4.2 Fake-News Detection
Figure 4: Node importance for fake-news spread graphs, over the GOS dataset. The node size correspond to its importance learned by the model, according to Equation 3
@@
Table 3: Evaluation of GMAN on the Gossipcop (GOS) fake news detection dataset.
@@
GMAN 97.34 ± 0.2
GossipCop (GOD) is a dataset of news articles annotated by professional journalists and fact-checkers, containing both content-based labels and social context information verified through the GossipCop fact-checking platform. It is composed of $5,464$ tree-structured graphs based on retweet information, where the news article is the root node and retweeting users are subsequent nodes in the cascade, with edges signifying retweet relationships. Each node in these graphs is associated with 4 features: 768-dimensional embeddings generated using a pretrained BERT model on user historical tweets, 300-dimensional embeddings from pretrained word2vec [27] on the same historical tweets, 10-dimensional features extracted from user profiles, and a 310-dimensional ”content” feature that combines the 300-dimensional embedding of user comments with the 10-dimensional profile features. Each graph is labelled to indicate whether it originates from a fake news post or not.
We decomposed the tree into a set of directed graphs rooted at the origin, each representing a distinct path of news propagation.
@@
We evaluated GMAN against $4$ GNNs, including GATv2 [5], GraphConv [29], GraphSAGE [18] and GCNFN [28]. We use random splits of train (80%), validation (10%), and test (10%) sets over the data and selected the best performing model over the validation set. We then report the average Accuracy score and std of the selected configuration with three seeds. Since many subgraphs reduce to a single node after decomposition, we group all size-one graphs into a shared subset and combine them non-linearly. For the remaining graphs, whose identities are not uniquely distinguishable, we apply a shared ExtGNAN. In total, we use two distinct ExtGNAN modules.
@@
Results are provided in Table 3. GMAN outperforms all baselines, with an uplift of 0.57 accuracy points.
Figure 4 presents the node importance of a random sample from the datasets, where the size of a node corresponds to its importance score according to Equation 3.

### v3

@@ 3.2 Node, graph and subset importance
$\text{TotalContribution}(j)=\sum_{l=1}^{K}[\mathbf{h}_{j}]_{F_{k}}=\sum_{w\in V}\rho\left(\Delta(w,j)\right)\sum_{l=1}^{K}\psi_{k}\left([\mathbf{x}_{w}]_{l},l\in F_{k}\right).$ (3)
@@
$\text{TotalContribution}(G)=\sum_{v\in G}\text{TotalContribution}(v).$
For graphs that are mixed non-linearly, i.e., graphs that belong in subsets of size greater than one, we provide instead the total contribution of the set to the final prediction
$\text{TotalContribution}(S)=\sum_{l=1}^{K}[\mathbf{S}]_{F_{k}}.$ (4)
@@ 3.3 Expressivity properties
In this section, we provide a theoretical analysis of the expressiveness of SuperMAN. Proofs are provided in the Appendix
@@
SuperMAN is strictly more expressive than GNAN.
The following theorem shows that a SuperMAN which is applied to subsets of graphs of size at least two, is more expressive than a SuperMAN that is applied to only subsets of size one:
@@
Let $S$ be a set of graphs $\{G_{i}\}_{j=1}^{m}$. Let $S_{1}=\{S_{i}\}_{i=1}^{m}$ be a partition of $S$ such that $|S_{i}|=1$. Let $S_{2}=\{S_{i}\}_{i=1}^{k}$ such that there exists $k$ with $|S_{k}|>1$. with a subset partition $\{S_{i}\}_{i=1}^{k}$. Then a SuperMAN trained over $S_{2}$ is strictly more expressive than a SuperMAN trained over $S_{1}$.
@@ 4 Empirical evaluation
In this section, we evaluate SuperMAN on real-world tasks, and demonstrate its interpretability properties  [^1] Code is provided in Supplementary Materials.
@@ 4.1 Medical predictions
We evaluate SuperMAN on two high-impact clinical prediction tasks: LoS of intensive care patients and onset of CD. In both settings, each individual is represented as a set of time-stamped biomarker trajectories, where each trajectory forms a directed path graph with nodes corresponding to test results and edges encoding the time elapsed between measurements. This representation preserves the temporal structure of each biomarker independently while enabling joint reasoning across biomarkers during learning.
@@
Next, we describe the two medical datasets used in our evaluation.
P12 (ICU Length of Stay): The PhysioNet2012 (P12) dataset, introduced by Goldberger et al. (2000), contains records from 11,988 intensive-care unit (ICU) patients, following the exclusion of 12 samples deemed inappropriate according to the criteria in Horn et al. (2020). For each patient, longitudinal measurements from 36 physiological signals were recorded over the initial 48 hours of ICU admission. Additionally, each patient has a static profile comprising 9 features, including demographic and clinical attributes such as age and gender. The dataset is labelled for a binary classification task: predicting whether or not the total LoS in the ICU exceeded 72 hours. The dataset is highly imbalanced, with $\sim$93% positive samples. To balance the classes we perform batch minority class upsampling, following the example in Zhang et al. (2021).
Crohn’s Disease (CD Onset): The Danish health registries (DHR) are comprehensive, nationwide databases covering healthcare interactions for over 9.5 million individuals (Pedersen, 2011). A key resource is the Registry of Laboratory Results for Research (RLRR), which has collected laboratory test results from hospitals and general practitioners since 2015 (Arendt et al., 2020). We first identify 8,567 individuals in the DHR who are later diagnosed with CD and use them as our patient class. We then construct a control pool by randomly sampling individuals from the DHR and downsampling by age to match the expected frequency of blood tests in the prediagnostic period, reflecting that CD typically manifests in early adulthood ( 20–30 years). From this age-matched control pool, we sample 8,567 controls, yielding two balanced classes. For each person, we extracted temporal trajectories of 17 routinely measured biomarkers, reflecting key physiological processes. The complete list and descriptions of these biomarkers are provided in the Appendix. The task is binary classification: predicting future CD onset from pre-diagnostic medical histories.
Table 1: Evaluation of SuperMAN on two real-world medical tasks. The metric reported is the mean AUPRC with standard deviation, calculated over 3 random seeds.
Methods LoS in ICU CD onset
Transformer 96.06 ± 0.32 75.60 ± 0.52
Trans-mean 96.44 ± 0.17 75.96 ± 0.92
GRU-D 95.91 ± 2.10 83.36 ± 0.40
SeFT 95.89 ± 0.08 71.22 ± 2.30
mTAND 93.02 ± 1.04 83.17 ± 0.67
DGM2 97.00 ± 0.40 83.02 ± 0.56
MTGNN 96.20 ± 0.78 75.26 ± 3.04
RAINDROP 96.32 ± 0.13 82.60 ± 0.82
SuperMAN 97.41 ± 0.38 83.93 ± 0.27
@@
We compare SuperMAN to $8$ baselines from Zhang et al. (2021), spanning both sequential and graph-based models, including: Transformer (Vaswani et al., 2017), a vanilla self-attention model applied to irregular time series; Trans-mean, which combines a Transformer with mean imputation of missing values; GRU-D (Che et al., 2016), a gated recurrent model with decay terms that encode informative missingness; SeFT (Horn et al., 2020), which treats each record as a set of timestamped feature-value pairs and aggregates them with permutation-invariant encoders; mTAND (Shukla and Marlin, 2021), a multi-time attention architecture that outperforms a broad range of RNN- and ODE-based models on irregular data; DGM2 (Wu et al., 2021) and MTGNN  (Wu et al., 2020), graph-based methods originally proposed for multivariate time-series forecasting; and Raindrop (Zhang et al., 2021), a state-of-the-art graph model for sparse, irregular EHR time series. For the $P12$ we used the splits as in  (Zhang et al., 2021). For CD, we randomly split the data into train (80%), validation (10%), and test (10%) sets. We conducted a grid search by training on the training set and evaluating on the validation set. We then selected the best-performing model over the validation set and report results over the test set. We define each individual biomarker as a unique signal type and group them according to coarse physiological categories. Importantly, these clinically-guided groups are based on broad domain knowledge and do not rely on specialized expertise. We also examine data-driven grouping that does not rely on any priors, based on the insights presented in Subsection 4.1.1. Due to space limitations inTable 4, App. F. We tune the biomarker subsets over no grouping at all (i.e, full interpretability), and $5$ additional subset groupings motivated by public common clinical knowledge. To account for class imbalance, and as commonly done for these datasets, we report the average AUPRC score and standard-deviation of the selected configuration with $3$ random seeds. Additional details on the datasets, experimental setup, subset groupings, and hyperparameter configurations are provided in the Appendix. We also present Calibration and robustness evaluations in Appendix G.
@@ Results
The results in Table 1 show that SuperMAN achieves the highest AUPRC across both medical prediction tasks. On Length of Stay in ICU, it improves upon the best baseline with a relative uplift of about $0.41$ points, while on Crohn’s Disease onset, it outperforms the best baseline by $0.57$ points. Notably, these improvements are achieved while SuperMAN also provides interpretability properties, as shown in detail in the next subsection.
@@ 4.1.1 Clinical insights through interpretability
Figure 2: Node-level importances for two individuals from (a) the P12 ICU LoS and (b) CD onset datasets. Node size indicates the exact node (measurement) contribution to the prediction.
Beyond predictive performance, a central strength of SuperMAN is that interpretability is built into the model by design, rather than added post hoc. This allows for fine-grained, temporally resolved explanations that go beyond global feature importance, enabling users to understand how and when specific biomarkers influence the model’s predictions. In high-stakes domains such as healthcare, this level of transparency is crucial. Clinical decision-making often depends not only on the outcome of a prediction but on a clear understanding of the reasoning behind it. Models that can provide such insights are far more likely to be trusted, audited, and integrated into clinical workflows. To highlight the interpretability of SuperMAN, we conduct attribution analyses on both the CD and P12 datasets, examining how the model assigns importance across time and signals.
@@
In the SuperMAN framework, nodes represent individual signals within a biomarker trajectory. As such, highly influential nodes can highlight critical phases where specific signals most strongly impact the prediction. We use Equation˜3 to quantify node-level contributions across biomarkers in both the CD and P12 datasets. To comply with privacy and data protection legislation, CD data is anonymised via noise and temporal shifting. Figure˜2 displays node-level importance across biomarker trajectories for two randomly selected individuals from each dataset, with node size reflecting the magnitude of each node’s contribution to the model’s prediction. In both clinical tasks, the model’s attributions appear consistent with established biomedical knowledge. For CD prediction, SuperMAN highlights key inflammatory and immune markers, such as F-Cal, platelets, and lymphocytes, as primary contributors. All of these are known to play central roles in disease onset (Vestergaard et al., 2023). In the P12 example, the model assigns high importance to markers of renal function, liver injury, cardiac stress, and metabolic imbalance, aligning well with clinical predictors of severity in intensive care settings.
@@
In addition to node-level importance scores, SuperMAN also provides subset-level scores, quantifying the contribution of entire biomarker groups to the model’s prediction. This enables flexible, system-level interpretability, allowing users to assess the collective influence of physiologically related biomarkers on the target variable. We conduct a subset-level importance analysis on the CD dataset using two clinically motivated grouping strategies. In the first, we assess individual biomarkers to determine whether the model prioritizes features known to be associated with CD onset, allowing us to verify its alignment with established biomedical knowledge. In the second, we group biomarkers into clinically coherent subsets based on physiological function, such as immune response; inflammation; oxygen transport; and liver function; to examine whether the model captures system-level patterns consistent with disease progression. Full details on the clinical significance of these groupings are provided in the Appendix.
(a) Single-biomarker groups SuperMAN
(b) Best-performing SuperMAN groups
@@
Figure 3: Subset-level contribution curves for Crohn’s Disease prediction. Each curve shows how the SuperMAN’s output changes as increasing noise is added to the latent representation of a biomarker group. (a) uses individual biomarkers; (b) uses physiologically coherent groups.
For each grouping strategy, we trained a separate instance of SuperMAN. To quantify the importance of each biomarker subset, we use Equation˜4, measuring how the model’s prediction changes when increasing the noise added to the features of nodes within that subset. To introduce noise in a structured way, we apply PCA (Abdi and Williams, 2010) to the feature vectors of all nodes corresponding to biomarkers in the group, and progressively perturb the input along the 1st principal component. This procedure is performed independently for each subset and allows us to assess the model’s sensitivity to perturbations in biologically meaningful groupings, offering insight into the relative predictive weight of each group. Importantly, SuperMAN is interpretable by design, which makes perturbation analysis directly reflect its internal computation. Unlike post-hoc attribution, we do not estimate influence indirectly but observe how the additive contribution changes. Adding structured noise (via PCA) and measuring the resulting prediction shift faithfully traces the model’s internal mechanism rather than relying on an external approximation. Thus, perturbation effects align exactly with feature importance, making the experiment both natural and principled. Figure˜3 presents the results of the subset-level attribution analysis. In the single-biomarker setting (Figure˜3(a)), F-Cal, platelets, and lymphocytes show strong directional effects on model output. F-Cal and platelets are positively associated with CD risk, while lymphocytes have an inverse effect, findings that are both biologically grounded and consistent with prior work (Vestergaard et al., 2023). In the clinically-coherent group setting (Figure˜3(b)), the inflammation subset emerges as the most influential, with a pronounced non-linear effect on predictions, aligning with established diagnostic relevance in CD (Vestergaard et al., 2023). We note that the above also serves as a quantitative measure of faithfulness, as in SuperMAN, interpretability is not post-hoc but built into the model architecture by design. Specifically, the contributions of nodes, graphs, or subsets are explicitly and additively used in computing the final prediction. This means that importance scores correspond directly to the actual values that are summed to produce the output label, making them faithful by construction.
@@ 4.2 Fake-News Detection
Table 2: Evaluation of SuperMAN on the Gossipcop (GOS) fake news detection dataset.
@@
SuperMAN 97.34 ± 0.2
GossipCop (GOS) is a dataset of news articles annotated by professional journalists and fact-checkers, containing both content-based labels and social context information verified through the GossipCop fact-checking platform. It is composed of 5,464 tree-structured graphs based on sharing information, where the news article is the root node and sharing users are subsequent nodes in the cascade, with edges signifying sharing relationships. Each node in these graphs is associated with 4 features: 768-dimensional embeddings generated using a pretrained BERT model on user historical posts, 300-dimensional embeddings from a pretrained word2vec (Mikolov et al., 2013) on the same historical posts, 10-dimensional features extracted from user profiles, and a 310-dimensional ”content” feature that combines the 300-dimensional embedding of user comments with the 10-dimensional profile features. Each graph is labelled to indicate whether it originates from a fake news post or not. We decomposed the tree into a set of directed graphs rooted at the origin, each representing a distinct path of news propagation.
@@
While SuperMAN can be applied to sets of graphs with any structure, the baselines used in the CD and P12 experiments are limited to path-like graphs and are unable to act on more complex graphs as in this dataset. Therefore, we instead evaluate SuperMAN against $4$ GNNs, including GATv2 (Brody et al., 2022), GraphConv (Morris et al., 2021), GraphSAGE (Hamilton et al., 2018) and GCNFN (Monti et al., 2019). We use random splits of train (80%), validation (10%), and test (10%) sets over the data and selected the best performing model over the validation set. We then report the average Accuracy score and std of the selected configuration with three seeds. Since many subgraphs reduce to a single node after decomposition, we group all size-one graphs into a shared subset and combine them non-linearly. For the remaining graphs, whose identities are not uniquely distinguishable, we apply a shared ExtGNAN. In total, we use two distinct ExtGNAN instances for this experiment.
Figure 4: Node importance for fake-news spread graphs, over the GOS dataset. The node size corresponds to its importance learned by SuperMAN, according to Equation˜3. All graphs with a single node are grouped into one subset. Therefore, the importance is provided on the subset level rather than the node level (green node).
@@
Results are provided in Table˜2. SuperMAN outperforms all baselines, with an uplift of 0.57 accuracy points. Figure˜4 presents the node importance of a random sample from the datasets, where the size of a node corresponds to its importance score according to Equation˜3.
4.2.1 Ablation study
Table 3: Ablation study of SuperMAN components.
Ablation AUPRC drop
[…]
(ii) $\rho\rightarrow$1 - 12.39% $\pm$ 1.39%
[…]
Finally, we carry out an ablation study on the CD dataset, to isolate the impact of key components on performance. Specifically, we take the best-performing configuration from the training-data grid search and re-train it with individual components ablated to assess their effect on performance. We test: (i) replacing DeepSet with mean pooling to assess the value of learned non-linear aggregation; (ii) replacing the distance function NN ($\rho$) with a constant 1 value.
This tests the usefulness of structural information; (iii) substituting ExtGNAN with a node-wise MLP to test the importance of graph inductive bias; (iv) replacing ExtGNAN with an identity mapping as a lower bound without feature learning; and (v) using standard GNAN (no multivariate feature groups) to evaluate the benefit of grouped feature processing. We report the performance degradation (AUPRC difference) in Table 3. The ablation results demonstrate that the core components of SuperMAN are critical to its effectiveness, as their removal consistently leads to notable performance degradation.